# font=small

### v1

@@ Abstract
In this paper, we study the effects of fill probabilities and adverse fills on the trading strategy simulation process. We specifically focus on a stochastic optimal control market-making problem and test the strategy on ES (E-mini S&P 500), NQ (E-mini Nasdaq 100), CL (Crude Oil) and ZN (10-Year Treasury Note), which are some of the most liquid futures contract listed on the CME (Chicago Mercantile Exchange). We provide empirical evidence which shows how fill probabilities and adverse fills can significantly effect performance, and propose a more prudent simulation framework for dealing with this. Many previous works aim to measure different types of adverse selection in the limit order book, however, they often simulate price processes and market orders independently. This has the ability to largely inflate the performance of a short-term style trading strategy. Our studies show that using more realistic fill probabilities, and tracking adverse fills, in the strategy simulation process, more accurately portrays how these types of trading strategies would perform in reality.
@@ 1 Introduction
Over the last 10-15 years, algorithmic trading and high-frequency trading (HFT) has become the most common method for completing trade transactions in the worlds leading financial markets. Trading algorithms can place many different types of trade orders, but this paper will solely focus on a strategy which uses market orders (MO) and limit orders (LO) to simplify the strategy simulation process. MOs are trade orders executed at the best available prices and LOs are trade orders which rest in the limit order book (LOB) until cancelled or executed by a corresponding MO. See Figure 1 for a visual description of the LOB in the E-mini S&P 500 contract (ES), where we selected an arbitrary time during the most active US stock market trading hours (9:30 EST - 16:00 EST). We will refer to these futures contracts by their ticker symbols from now on.
Mathematical finance experts have long used stochastic processes to model the variables in algorithmic and HFT markets, where Cartea et al., (2015) brings together a whole range of Stochastic Optimal Control (SOC) examples with many variations of trading problems related to liquidation, acquisition, market making, volume-targeted trading, pairs trading, statistical arbitrage and order imbalance. The type of problem that this paper will focus on is the Market-Making (MM) style problem given in section 10.4.2 of Cartea et al., (2015), where the stochastic control is whether or not the market-maker should be posted on the best bid/ask. This problem was chosen as it highlights two of the major issues in MM that are often overlooked and oversimplified in the literature, the fill probability and the adverse fill. These can be summarized as follows:
  • Fill Probability: Measures the probability of a traders LO being filled when they’re posted in the LOB. This often depends on the depth in the LOB and the average trade size. As can be seen in Figure 1, one measure of the depth is the size of the bid and ask queues displayed in the LOB. One could also look at the volume traded relative to these LO sizes and in Table 1 we show the mean and median trade size per contract in four of the most liquid US futures contract, where we used LOB data from April 9th 2024 to calculate this.
@@
  • Adverse Fills: A form of adverse selection which occurs when a passive MMs LO is picked off and they receive disadvantageous prices on their filled trade orders. In other words, immediately after their trade order is filled, their new trade position is out of the money when marked-to-market. We will show how the adverse fill occurs on the majority of trade order fills across many different assets, and based on this empirical evidence, we devised a method for including these adverse fills in our proposed trading simulation environment.
@@
This paper will proceed as follows. We begin by discussing the significance of the paper along with a motivating example. In section 2, we will go over some of the literature on fill probabilities and adverse fills. In section 3, we introduce a SOC MM problem, which we will use for our simulation analysis. In section 4, we explain how one can create a simulation environment by discretizing the continuous processes given in section 3. Here we compare the simulation environment from Cartea et al., (2015) and Jaimungal, (2019), which we will refer to as the benchmark environment, with our improved simulation environment. Then, in section 5, we show the results of the performance of the trading strategy in both simulation environments. Lastly, we finish this paper off with some concluding remarks and ideas for future research.
1.1 Motivation and Significance of the Paper
The goal of this paper will be to contribute to the existing literature on the trading strategy simulation process by portraying a more accurate way to simulate performance. Most trading simulation frameworks in the literature model price processes and market orders independently, which can often lead to significantly inflated performance measures. The first problem here is that adverse fills often end up being excluded from the trading strategy simulation process, and, secondly, unrealistic fill probabilities are often used for every other type of trade order fill, which we will refer to as non-adverse fills. To stress the importance of tracking adverse fills in the simulation process, one must recall that in a real live trading environment, if a LO is posted at a specific price, and the price of the asset ends up going through this price, you are essentially guaranteed to have gotten a fill at a worse price than the current price. This is merely based on the rules of the LOB system. Thus, excluding any trade order fills of this nature, from the trading strategy simulation process, would create an unrealistic fill tracking process.
@@ 1.1 Motivation and Significance of the Paper
To go into a bit more detail on the importance of this, we give an example by analysing some limit order fill data from a professional simulator environment, using the Trading Technologies (TT) software platform. The TT Simulation environment is a SaaS technology trading platform developed by a company called Trading Technologies. As well as enabling users to trade in live markets, TT has developed its own simulation environment where users can test strategies without risking real capital. Simulation performance results here are often highly in-line with what would happen in reality, although it is important to point out that a major exception would be for strategies that aim to trade high relative volume at specific prices. Using the TT platform, we will give some empirical evidence on how often adverse and non-adverse fills occur based on a sample of simulated trades throughout a given trading day. TT can quite accurately simulate when you are likely to receive a LO fill. Put simply, whenever a LO is placed, it is placed in the LOB and given an imaginary queue position which would be the same as what one would have been given in reality. Then, a LO is considered filled when the trade order behind the imaginary order in the queue is filled in reality. This method can simulate fills quite realistically, particularly if the order size is relatively small, as this traders flow is unlikely to have effected prices in reality.
@@
To show the likelihood of adverse fills occurring, we start by simulating a simple trading strategy in ES, NQ, CL and ZN, which posts limit orders close to the best bid/ask over a single trading day. For ES and CL, the strategy placed static orders on the bid and ask 4 ticks away from each other, in NQ 16 ticks away, and in ZN 1 tick away, while also placing orders at the same filled price again once a fill occurs at a different price. We picked these levels so that the basic posting strategy would trade low relative volume throughout the day, but at the same time, complete enough trades so that we can study the empirical limit order fill data. Each fill that precedes a disadvantageous move in the price will be considered adverse, while every other fill will be considered non-adverse. More specifically, when a LO is filled on the bid (ask) and the next new bid (ask) price is lower (higher), we consider this fill to be adverse, as the first change in the marked-to-market value of the position is negative. By marked-to-market, we used the price that the position can immediately be liquidated at using a MO, which involves crossing the spread. In order to check this, we iterated over each future LOB event until the first change in the marked-to-market value for each individual trade was nonzero. See Table 2 for the results, where we can see that a significant portion of the total number of LO fills in ES, NQ, CL and ZN were adverse.
@@
Thus, based on the above empirical results, it is quite obvious that adverse fills could significantly effect the performance of any short-term intra-day trading strategy posting LOs, which leads us to believe that a trading simulation environment should account for this. Later on, we develop an improved simulation environment in a MM setting, where we show that the results significantly change when the simulation process tracks adverse fills, as well as utilizing non-adverse fill probabilities. We chose to perform our analysis in a MM setting, since MM strategies often trade passively, in the sense that MMs often post a large amount LO’s in the LOB and aim to trade a large amount volume throughout any given day.
2 Literature Review
There are some recent papers in optimal MM that try to incorporate the effects of fill probabilities and adverse selection into the LOB system and the trading strategy process, while there also some that use SOC theory to setup the trading problem. For example, Bulthuis et al., (2017) devises a model for analyzing LOs and MOs and how they create a permanent and temporary price impact, while also including a component for unfilled LOs, under a SOC framework. Cartea et al., 2018a developed a volume imbalance indicator to make predictions about future MOs. They designed a Markovian pure jump model of the price process, under a SOC framework, including variables such as the spread, LO and MO arrivals and volume imbalance. Roldan Contreras and Swishchuk, (2022) study various types of SOC trading problems, including MM, where they aim to portray some of the unique characteristics on how asset prices move through a dependence on recent activity and jump features. Their optimal solutions are given in terms of the arrival rates of trade orders, which they model utilizing the General Compound Hawkes Process framework. Cartea et al., 2018b , Cartea et al., (2015) and Cartea et al., (2014) develop models which assumes HFT has superior information where they can utilize this to post and cancel LOs at a quicker rate. Their models analyze factors such as arriving MOs, trade clustering and variables that can lead to a change in the shape of the LOB. They create a short-term alpha predictor with the hope of predicting future adverse fills so that they can avoid them. While collecting superior information like this may help in certain instances when making future price predictions, it cannot avoid regular random day-to-day trade order flow where it is normally too late to change the order posting strategy when random spikes in activity instantly enter the market. Predictions, in general, can often be wrong too. The instant nature of trading activity means it is often too late to change the LO posting strategy when random spikes occur. We will discuss this in more detail later from section 4 onward.
From a non SOC standpoint, there is also quite a bit of research on LOBs and fill probabilities. For a broad survey on how LOBs are modelled, see Gould et al., (2013). They discuss a large portion of the different methods used to model LOBs from a wide array of empirical and theoretical research studies. They also highlight the many shortcomings and obvious empirical facts that are yet to be included within these models. Cont and De Larrard, (2013) devised a stochastic framework for modelling the dynamics of the LOB where they made computations based on the arrival of MOs and LOs through a Markovian queueing system. Their setting made it easy to compute numerous important variables such as the distribution for the time length between changes in price, the distribution and autocorrelation for the changes in price and predictions for midprice increases/decreases. In Swishchuk and Vadori, (2017) they extend their framework by creating an arbitrary distribution for the inter-arrival times of LOB events, as well as creating a dependence structure between new and previous LOB book events, under a Markov Renewal process framework. In Swishchuk et al., (2019), they introduce two types of Hawkes processes: compound and regime-switching. Through these models they express price volatility in terms of parameters describing arrival rates and price changes. Maglaras et al., (2022) uses a recurrent neural network to predict the fill probabilities of LOs. It is the first study that directly predicts LO executions via the distribution of fill times. Arroyo et al., (2024) develops a model using survival analysis which maps time-varying features of the LOB to the distribution of fill times of limit orders. Their method uses a convolutional-transformer encoder and a monotonic neural network decoder. Lokin and Yu, (2024) proposes a queueing system method for calculating the fill probabilites of LOs posted at different levels in the LOB. Their probability model covers a wide array of scenarios including midprice changes, probabilities for orders posted at the best bid/ask and at deeper levels in the LOB. Hoffmann, (2014) devises a framework whereby you have fast and slow traders interacting together, where the fast traders can quickly revise their quotes after new information enters the market, which thus creates the ability to avoid being picked off. Cont et al., (2014) examines the impact of LOB events and that price movements occur mostly due to order flow imbalance. They portray how this could be used as a measure of adverse selection in LOs. Easley et al., (2012) discusses a method for measuring how toxic order flow can effect the MM adversely, based on volume imbalance and intensity of orders. They argue it is a good model for predicting toxic order flow over a short-term horizon. Brogaard et al., (2019) devises a model for analyzing how MOs and LOs entered by HFT and non-HFT traders induce price discovery, which they argue mostly comes from LOs, and the majority of these orders come from HFTs. They portray how a decrease in the volume of LOs entering the market often leads to increased volatility in the price.
As described above, some forms of adverse selection in relation to adverse fills have been incorporated into different frameworks in the literature. However, in terms of correcting the trading process or the LOB models for adverse fills, we have seen no prior work. No matter how good a proposed trading model is, it will be virtually impossible to avoid all adverse fills. DeLise, (2024) is the first paper to give an in-depth analysis of the likelihood of adverse fills, where they show some empirical evidence for how often adverse fills occur. They portray how the majority of trade order fills from the financial asset they study, ZN, are, in fact, adverse. We extended their ideas by also analysing the rate of adverse fills in ZN in Table 2, as well as some additional products, to further justify the empirical evidence on adverse fills, which made it clear it’s paramount to include these in any short-term trading strategy evaluation process.

### v2

@@ Abstract
In this paper, we study the effects of fill probabilities and adverse fills on the trading strategy simulation process. We specifically focus on a stochastic optimal control market-making problem and test the strategy on ES (E-mini S&P 500), NQ (E-mini Nasdaq 100), CL (Crude Oil) and ZN (10-Year Treasury Note), which are some of the most liquid futures contracts listed on the CME (Chicago Mercantile Exchange). We provide empirical evidence that shows how fill probabilities and adverse fills can significantly affect performance and propose a more prudent simulation framework to deal with this. Many previous works aim to measure different types of adverse selection in the limit order book (LOB), however, they often simulate price processes and market orders independently. This has the ability to largely inflate the performance of a short-term style trading strategy. Our studies show that using more realistic fill probabilities and tracking adverse fills in the strategy simulation process more accurately shows how these types of trading strategies would perform in reality.
@@ 1 Introduction
Over the last 10-15 years, algorithmic trading and high-frequency trading (HFT) has become the most common method for completing trade transactions in the worlds leading financial markets. Trading algorithms can place many different types of trade orders, but this paper will focus solely on a strategy that uses market orders (MO) and limit orders (LO) to simplify the strategy simulation process. MOs are trade orders executed at the best available prices, and LOs are trade orders that rest in the limit order book (LOB) until canceled or executed by a corresponding MO. See Figure 1 for a visual description of the LOB in the E-mini S&P 500 contract (ES), where we selected an arbitrary time during the most active trading hours (9:30 EST - 16:00 EST) of the US stock market. We will refer to this futures contract by its ticker symbols from now on.
Mathematical finance experts have long used stochastic processes to model variables in algorithmic and HFT markets, where Cartea et al., (2015) brings together a wide range of Stochastic Optimal Control (SOC) examples with many variations of trading problems related to liquidation, acquisition, market making, volume-targeted trading, pairs trading, statistical arbitrage, and order imbalance. The type of problem that this paper will focus on is the Market Maker (MM) style problem given in Section 10.4.2 of Cartea et al., (2015), where the stochastic control is whether or not the MM should be post one unit (be it shares/lots etc) on the best bid/ask. This problem was chosen because it highlights two of the major issues for a MM that are often overlooked and oversimplified in the literature, the fill probability and the adverse fill. These can be summarized as follows:
  • Fill Probability: The likelihood that a trader’s LO will be executed after being posted in the LOB. This probability is influenced by factors such as the position of the order in the queue, the depth of liquidity at that price level, and the frequency and size of incoming market orders. As shown in Figure 1, one way to measure depth is by examining the size of the bid and ask queues in the LOB. Another approach is to analyze the ratio of the traded volume to the resting LO sizes. Table 1 provides the mean and median trade sizes per contract for four of the most liquid U.S. futures contracts, calculated using LOB data from April 9, 2024.
  • Adverse fills: A form of adverse selection that occurs when a passive MMs LO is executed at a disadvantageous price. Specifically, this happens when the MM’s order is “picked off,” which means that immediately after execution, the new trade position is out of the money when marked-to-market. We will demonstrate that adverse fills occur in the majority of trade executions across various assets. Based on this empirical evidence, we develop a method to incorporate adverse fills into our proposed trading simulation environment, ensuring a more realistic representation of execution risk.
@@
The main contributions of the paper can be summarized as follows:
  • We propose an enhanced simulation framework for evaluating the performance of HFT strategies, particularly in market-making. This framework accounts for both adverse and non-adverse trade order fills in a more realistic manner, addressing a key limitation in many existing approaches that assume independence between trade order fills and price processes. Our framework highlights how this assumption can significantly affect performance.
  • We introduce a probability measure for non-adverse trade order fills, refining how these fills are modeled within MM strategies. As will be demonstrated in Section 3, simplistic strategies often suffer from excessive adverse fills and are not guaranteed to receive non-adverse fills for every incoming market order. While fill probabilities are not new, they should only be applied to non-adverse fills, as adverse fills are guaranteed to occur based on the rules of most LOB systems, a limitation we address here. A more realistic probability measure ensures that performance metrics align better with real-world trading conditions.
  • We analyze a well-studied MM problem using the framework of Cartea et al., (2015), extending their SOC formulation to incorporate non-adverse fill probabilities in the optimal solutions. Through an empirical study on four highly liquid CME futures contracts, we demonstrate the significant impact of both non-adverse and adverse fills on performance. Although adverse fills could not be directly incorporated into the SOC framework due to necessary dimension reduction techniques, our simulation environment still allows us to assess their implications on performance.
@@
The rest of this paper will proceed as follows. In Section 2, we will go over some of the literature on fill probabilities and adverse fills. In Section 3, we begin by discussing the significance of the paper with two simple motivating examples of a MM trading strategy. In Section 4, we introduce an SOC MM problem, which we will use for our simulation analysis. In Section 5, we explain how one can create a simulation environment by discretizing the continuous processes given in Section 4. Here we compare the simulation environment in Cartea et al., (2015) and Jaimungal, (2019), which we will refer to as the benchmark environment, with our improved simulation environment. Then, in Section 6, we show the performance results of the trading strategy in both simulation environments. Lastly, this paper ends with some concluding remarks and ideas for future research.
2 Literature Review
Recent research on optimal MM problems incorporates fill probabilities and adverse selection in the LOB system and the trading strategy process, with some approaches leveraging SOC theory. For example, Bulthuis et al., (2017) develops a model analyzing the price impact of LOs and MOs, incorporating a component for unfilled LOs within an SOC framework. Cartea et al., 2018a introduces a volume imbalance indicator to predict future MOs, using a Markovian pure jump model that accounts for the spread, LO/MO arrivals, and volume imbalance. Roldan Contreras and Swishchuk, (2022) explores various SOC trading problems, including MM, highlighting how asset prices evolve based on recent activity and jump characteristics. Their optimal solutions are expressed in terms of trade order arrival rates, modeled via the General Compound Hawkes Process framework. Furthermore, Cartea et al., 2018b , Cartea et al., (2015), and Cartea et al., (2014) examine HFT strategies that aim to leverage superior information to post and cancel LOs more efficiently. Their models analyze MO arrivals, trade clustering, and LOB shape dynamics, using a short-term alpha predictor to anticipate and avoid adverse fills. However, while such predictive models may offer advantages in some scenarios, they cannot fully account for random, HFT order flow, where sudden spikes in activity occur too quickly for an effective response. Predictions can also be unreliable, especially when the short-term alpha predictor lacks a meaningful market edge. We explore this issue further in Section 4, where we analyze a similar MM problem.
Beyond SOC models, there is extensive research on LOBs and fill probabilities. For a broad survey, Gould et al., (2013) reviews empirical and theoretical modeling approaches, highlighting key shortcomings. Cont and De Larrard, (2013) introduces a stochastic LOB framework using a Markovian queuing system to model MO and LO arrivals, enabling computations of price change distributions, autocorrelations, and midprice movements. Swishchuk and Vadori, (2017) extends this by allowing arbitrary inter-arrival distributions and dependencies between LOB events under a Markov Renewal process framework. Swishchuk et al., (2019) further enhances this by incorporating compound and regime-switching Hawkes processes, expressing price volatility in terms of arrival rates and price changes. Maglaras et al., (2022) employs a recurrent neural network to predict LO fill probabilities, marking the first study to model executions via fill time distributions. Arroyo et al., (2024) applies survival analysis with a convolutional transformer encoder and a monotonic neural network decoder to map LOB characteristics to fill time distributions. Lokin and Yu, (2024) develops a queuing system for calculating LO fill probabilities at different depths, considering midprice changes and best bid/ask orders. Hoffmann, (2014) models interactions between fast and slow traders, where fast traders update quotes rapidly to avoid being picked off. Cont et al., (2014) links price movements to order flow imbalance, using it as a measure of adverse selection. Easley et al., (2012) proposes a model linking toxic order flow to MM risk via volume imbalance and order intensity, offering a short-term predictive framework. Brogaard et al., (2019) analyzes how MO and LO submissions by HFT and non-HFT traders drive price discovery, finding that most price information stems from LOs, particularly from HFTs. They also show that reduced LO volume leads to increased price volatility.
Although some frameworks incorporate adverse fills, no prior work has explicitly adjusted the trading process or LOB models to distinguish between adverse and non-adverse fills. Regardless of the quality of the model, it is virtually impossible to completely avoid adverse fills. DeLise, (2024) is the first to analyze the likelihood of adverse fills, providing empirical evidence on their frequency. They show that in ZN, most trade order fills are adverse. We extend their work by examining the adverse fill rate in ZN (Table 2) and additional products, further demonstrating the necessity of incorporating adverse fills in short-term trading strategy evaluations. In Section 3, we introduce a method to distinguish between adverse and non-adverse fills through simple MM examples. In Section 4, we present a more complex SOC MM problem, and in Section 5, we propose a method to integrate this trade order fill logic into the framework.
3 Adverse and Non-adverse Trade Order Fills
Most existing simulation models treat price processes and market orders independently, often inflating performance metrics. The primary issue with this is that adverse fills are frequently excluded, while unrealistic fill probabilities are applied to other trade executions, which we refer to as non-adverse fills. Accurately tracking adverse fills is crucial because in real trading environments, if a LO is posted at a given price and the asset price moves through it, the order is effectively guaranteed to be filled at a worse price. This is a fundamental property of the LOB system. Excluding such fills leads to an unrealistic fill-tracking process. Similar findings to ours also appear in Law and Viens, (2019) and DeLise, (2024), where the former highlights how diffusion-based market-making models can create “phantom gains”, and the latter provides an empirical analysis of the “negative drift of a limit order fill”, a key factor in an MM’s performance.
To demonstrate this, consider a real-world LOB observed at 1-second intervals, as we will later in our empirical analysis. In most assets, the best bid and ask prices remain unchanged significantly more often than they move at each time step. This is supported by empirical analyses in several seminal works, including Biais et al., (1995), Hasbrouck and Saar, (2013), and Cont et al., (2014), among others. If market orders are modeled independently of price changes, a trade order fill does not necessarily align with the price moving through an order. Unless the model explicitly enforces fills when price levels are crossed, randomly generated executions are unlikely to capture adverse fills, systematically underestimating their frequency. This issue worsens as the time step decreases. With fewer price changes per step, fills become even less likely to align with price movements, distorting performance assessments, particularly in HFT settings. We will also give empirical evidence in Section 5 that demonstrates how failing to account for adverse fills leads to a significant misrepresentation of strategy performance.
In the remainder of this section, we present two examples to illustrate key concepts and results. Section 3.1 introduces a simple example to clarify the distinction between adverse and non-adverse fills, serving purely as an illustration of the two trade-order fill types. Section 3.2 explores a basic LO posting strategy for an MM, simulated in a real-time trading environment. This second example provides insight into the expected distribution of adverse and non-adverse fills for a simple MM strategy, particularly in the context of relatively small order sizes.
3.1 Example 1: Simple Market Maker Trade Order Fills
Consider a simple example that simulates an MM scenario where an LO is always posted on the best bid and ask, and one MO arrives at each time step, randomly filling the MM agent on the bid or ask. The best bid and ask prices evolve as a random walk with a fixed 1-cent spread. Each fill is classified as adverse (if the next price move is unfavorable) or non-adverse (if the price remains the same or moves favorably). It is assumed that the MM agent’s LOs are always at the front of the queue and of size 1, thus, they are fully filled with every incoming MO. This can also be interpreted as the MM having a fill probability of 100%. See Figure 2, which visualizes bid and ask prices over time, with red markers for adverse fills and green markers for non-adverse fills, clearly illustrating the different trade order fill types. In this plot, the best bid and ask prices are represented by blue and yellow lines, respectively. Note the following:
  • The point of this example is merely to highlight the difference between adverse and non-adverse fills.
  • While the time steps are uniform in this example, they are non-uniform in reality.
Figure 2: A visualization of the simple MM strategy in Example 1, highlighting the timing of each trade order fill and distinguishing between the different trade order fill types.
3.2 Example 2: Market Maker Trade Order Fills in a Real-Time Simulation Environment
@@
To go into a bit more detail on the importance of distinguishing between adverse and non-adverse trade order fills, we give an example by analyzing some LO fill data from a professional simulator environment, using the Trading Technologies (TT) software platform. The TT Simulation environment is a trading platform developed by a company called Trading Technologies. In addition to providing access to live markets, TT has created a simulation environment that allows users to test strategies without exposing real capital to risk. Simulation performance results here are often highly in line with what would happen in reality, although it is important to point out that a major exception would be for strategies that aim to trade high relative volume at specific prices. Using the TT platform, we will give some empirical evidence on how often adverse and non-adverse fills occur based on a sample of simulated trades throughout a given trading day (from 8:30 EST-16:00 EST). TT can quite accurately simulate when you are likely to receive a LO fill. Simply put, when an LO is placed, it is assigned an imaginary queue position in the LOB, mirroring the position it would have received in a real market. The LO is then considered filled when a real trade order is executed directly behind its imaginary queue position. This approach provides a realistic fill simulation, especially for relatively small order sizes, as such trades are unlikely to impact market prices significantly.
To show the likelihood of adverse fills occurring, we start by simulating a simple trading strategy in ES, NQ, CL and ZN, which posts LOs close to the best bid/ask over a single trading day. The main components of the strategy can be described as follows:
  1. 1.
    For ES and CL, the strategy placed static orders for 1 lot at bid and ask prices 4 ticks away from each other, in NQ 16 ticks away, and in ZN 1 tick away. We picked these levels so that the basic posting strategy would trade low relative volume throughout the day, but at the same time, complete enough trades so that we can study the empirical LO fill data. The different tick size spacing in the posted LOs is related to how actively each of these assets trades, as well as how often price changes occur in the asset. For example, NQ experiences the most frequent price changes compared to ES, CL, and ZN, while ZN has the fewest and often remains unchanged for extended periods. These adjustments allowed us to obtain a reasonable and comparable amount of data for each asset.
  2. 2.
    The strategy also reposts orders at the same filled price whenever a fill occurs at a different price. A sample set of steps that this strategy could follow, where here the focus will be on CL, can be summarized as follows:
    1. i.
      At $t=1$, where $t$ represents a non-uniform timestep determined by event occurrences in the LOB, the MM agent posts two LOs in CL: a buy LO at $81.86$ and a sell LO at $81.90$. The price at $t=1$ lies somewhere between these two price levels.
    2. ii.
      At $t=2$, the buy LO is filled at $81.86$.
    3. iii.
      At $t=3$, the agent places a new buy LO at $81.82$ while keeping the existing sell LO at $81.90$ unchanged.
    4. iv.
      At $t=4$, the new buy LO is filled at $81.82$.
    5. v.
      At $t=5$, the agent places a new buy LO at $81.78$ and a new sell LO at $81.86$.
    6. vi.
      At $t=6$, the new sell LO is filled at $81.86$.
    7. vii.
      At $t=7$, the agent places a new buy LO at $81.82$ and, if not already posted, a new sell LO at $81.90$. In this example, previously posted orders that are not between the highest posted bid LO and the lowest posted ask LO remain active, though one could also consider canceling resting LOs further from the market.
    This process continues in a similar fashion throughout the trading day. The primary motivation for this LO posting sequence is to prevent the fill data from being skewed by a high concentration of trades occurring at the same price within a short time window. This ensures a more representative distribution of adverse and non-adverse fills while also avoiding repetitive oscillatory patterns that may arise over short time intervals.
@@
As in example 1, each trade order fill that precedes a disadvantageous move in the price will be considered adverse, while every other fill will be considered non-adverse. More specifically, when an LO is filled on the bid (ask) and the next new bid (ask) price is lower (higher), we consider this fill to be adverse, as the first change in the marked-to-market value of the position is negative. See Table 2 for the results, where we can see that a significant portion of the total number of LO fills in ES, NQ, CL and ZN were adverse.
@@
Thus, the empirical results above clearly demonstrate that adverse fills can significantly impact the performance of any short-term intraday trading strategy that posts LOs. This highlights the need for a trading simulation environment that properly accounts for adverse fills, particularly when analyzing a simple MM strategy. In Section 4, we introduce an SOC optimal MM problem, where this simple strategy is unlikely to predict future order flow with high accuracy or anticipate price movements as effectively as professional MMs, who typically employ more sophisticated strategies. Unfortunately (for academics), most of these strategies remain highly confidential, as revealing them would lead to alpha decay (see Zhou and Lin, (2017) and Sivaramakrishnan et al., (2018) for an empirical analysis of this phenomenon), making them difficult to replicate. Although replicating more sophisticated strategies is challenging, their performance assessment measures can still be studied and further improved, as these measures are largely strategy independent. Consequently, in our SOC MM scenario, we should expect a similar distribution of adverse and non-adverse fills.
Later, we develop an improved MM simulation environment, demonstrating that results change significantly when the model explicitly tracks adverse fills and incorporates realistic non-adverse fill probabilities. We focus on MM strategies because MMs typically trade passively, posting large volumes of LOs in the LOB and executing a substantial number of trades throughout the day.